# JL1-CD: A New Benchmark for Remote Sensing Change Detection and a Robust Multi-Teacher Knowledge Distillation Framework

### v1

@@ Abstract
Deep learning has achieved significant success in the field of remote sensing image change detection (CD), yet two major challenges remain: the scarcity of sub-meter, all-inclusive open-source CD datasets, and the difficulty of achieving consistent and satisfactory detection results across images with varying change areas. To address these issues, we introduce the JL1-CD dataset, which contains 5,000 pairs of 512 × 512 pixel images with a resolution of 0.5 to 0.75 meters. Additionally, we propose a multi-teacher knowledge distillation (MTKD) framework for CD. Experimental results on the JL1-CD and SYSU-CD datasets demonstrate that the MTKD framework significantly improves the performance of CD models with various network architectures and parameter sizes, achieving new state-of-the-art results. The code is available at https://github.com/circleLZY/MTKD-CD.
@@ I INTRODUCTION
In recent years, the rapid advancement of deep learning (DL) has revolutionized remote sensing CD, delivering substantial performance breakthroughs. DL-based CD methods generally involve three steps: (1) extracting change features from image pairs, (2) generating change maps based on the extracted features, and (3) predicting labels based on the feature maps. Convolutional Neural Networks (CNNs), which achieved remarkable success in image processing, were the first neural network architecture applied to remote sensing CD and remain widely optimized and utilized today [10, 11, 12]. With the introduction of Transformers, some studies have explored their application in CD tasks [13, 14]. More recently, Foundational Model (FM) has emerged as a novel paradigm, aiming to achieve multi-task and multi-domain generalization through large-scale pretraining [15]. There have been works utilizing remote sensing data to fine-tune pretrained models such as Vision Transformers (ViT)[16], Segment Anything Model (SAM)[17], and Contrastive Language-Image Pretraining (CLIP)[18], achieving higher performance in CD tasks [19, 20, 21, 22]. Several semi-supervised methods have also been proposed, further improving CD performance under limited labeled data. Zhang et al.[23] propose a novel multilevel consistency-regularization-based semi-supervised CD approach, incorporating Fourier-based frequency transformation and a dynamic pseudolabel selection scheme to mitigate background noise and improve unlabeled data utilization. Xu et al. [24] introduce a semi-supervised label and embedding consistency network (SS-LEC) for ORSI scene classification, which strategically enforces consistency across augmentations and stages of training. Li et al. [25] propose SemiCD-VL, a VLM-guided semi-supervised change detection method that synthesizes pseudo labels via a mixed change event generation strategy, achieving significant performance gains over FixMatch and SOTA unsupervised methods.
However, DL-based CD methods generally face two major challenges: the scarcity of high-quality, high-resolution, all-inclusive CD datasets and limitations in handling highly dynamic change areas. Although numerous CD datasets have been constructed and proposed, they are often tailored to specific scenarios, which restricts the generalization capabilities of the algorithms. For instance, models trained on datasets focused on human-induced changes often fail to perform effectively when confronted with natural change scenarios. On the other hand, the learning capacity of these models is inherently limited. Most existing algorithms rely on a single-phase training approach, typically end-to-end training, supplemented by techniques such as learning rate decay. While such training strategies can achieve satisfactory results within a constrained range of changes, the models’ performance significantly degrades when addressing scenarios with wide variations in change areas, ranging from no change to complete change.
Figure 1: Timeline of the development of mainstream DL-based CD methods.
@@ II-B DL-Based CD
Figure 2: Sample images from the JL1-CD dataset. Each row, from top to bottom, represents: the image at time 1, the image at time 2, and the ground truth label. Each column corresponds to different change types: (a) Decrease in woodland; (b) Building changes; (c) Conversion of cropland to greenhouses; (d) Road changes; (e) Waterbody changes; and (f) Surface hardening (central region).
@@
Recently, foundation models have become a new training paradigm, and some works have applied them to CD tasks. Li et al. [19] proposed the Bi-Temporal Adapter Network (BAN), a universal FM-based framework for CD, which enhances existing models with minimal additional parameters and achieves significant performance improvements. Chen et al. [21] introduced Time Travelling Pixels (TTP), a method that integrates latent knowledge from the SAM model into CD, overcoming domain shifts and spatio-temporal complexities, demonstrating SOTA results on the LEVIR-CD[32] dataset. Zheng et al. [22] developed AnyChange, a zero-shot CD model built on the SAM that utilizes bitemporal latent matching for training-free adaptation, setting a new SOTA on the SECOND[39] benchmark and achieving significant improvements in both unsupervised and supervised CD tasks.
@@ III JL1-CD Dataset
We collect the number of types of changes, number of image pairs, image size, and resolution information of mainstream CD datasets in Table I. The SZTAKI AirChange Benchmark [46] contains 12 pairs of 952 × 640 and one pair of 1,048 × 724 optical aerial images. It is one of the earliest and most commonly used CD datasets in early research. The DSIFN dataset [31] consists of 6 large bi-temporal image pairs from 6 cities in China, which are cropped into 394 sub-image pairs, each sized 512 × 512. The SECOND dataset [39] includes 4,662 pairs of aerial images collected from multiple platforms and sensors, covering cities such as Hangzhou, Chengdu, and Shanghai. Unlike other datasets that classify changes into only two categories (change and no change), SECOND provides detailed annotations for change types, including non-vegetated surfaces, trees, low vegetation, water, buildings, and playgrounds. However, the resolution of all or part of the images in these datasets does not reach the sub-meter level. WHU-CD [47], LEVIR-CD [32], and S2Looking [48] are very popular datasets that are specifically designed for monitoring building changes. These datasets predominantly include human-induced changes and lack natural change types. The CDD dataset is derived from 7 pairs of 4,725 × 2,700 real-world seasonal change remote sensing images. The SYSU-CD contains 20,000 pairs of 0.5-meter aerial images captured in Hong Kong between 2007 and 2014. These datasets feature high resolution and diverse change types.
To provide a better benchmark for evaluating CD algorithms, we propose the JL1-CD dataset, a high-resolution, all-inclusive change detection dataset. JL1-CD includes 5,000 pairs of satellite images captured in China from early 2022 to the end of 2023, including Shandong, Ningxia, Anhui, Hebei, Hunan, and other regions. The images have sub-meter resolutions ranging from 0.5 to 0.75 meters and are sized 512 × 512 pixels. As shown in Fig. 2, the dataset covers various common human-induced and natural surface features, such as piled earth, buildings, roads, hardened surfaces, woodlands, grasslands, croplands, and water bodies. The original 5,000 image pairs are divided into 4,000 pairs for training and 1,000 pairs for testing, following an 80:20 split. The JL1-CD dataset will be made openly available for all research needs.
@@ IV-A O-P Strategy
The traditional training and testing approach for CD models is illustrated in the upper-left corner of Fig. 3 (green box) and the upper-right corner (pink box). For a given change detection model $\mathcal{M}$, the input consists of a pair of images $(X_{1},X_{2})$, and the output is a change map (CM). If the number of channels $c=1$, the predicted label $\hat{Y}$ can be directly obtained using a thresholding method as follows:
@@
where $(i)$ denotes the pixel location in the image, ”1” represents a change, and ”0” represents no change (for visualization purpose, ”1” will be mapped to a grayscale value of 255). The threshold $th$ is a predefined value. If $c=2$, the predicted label $\hat{Y}$ is typically determined using an element-wise comparison of the two channels, such as:
@@
where $CM_{0}$ and $CM_{1}$ represent the first and second layers of the change map, respectively. If $c>2$, it indicates that the CD model $\mathcal{M}$ not only predicts the locations of changes but also detects the types of changes, which is beyond the scope of this paper.
@@ IV-B MTKD Framework
By partitioning the training set, we effectively reduce the learning burden for each model. As a result, the Origin-Partition (O-P) strategy significantly enhances the performance of the CD algorithm. However, a clear issue arises: during inference, we are required to load 4 different models, and even disregarding data throughput, the time spent is at least twice that of the original algorithm (often much more). This substantially increases both memory and time complexity. We are thus prompted to consider: is there a way to combine the capabilities of these four models into a single model?
To address this, we propose the Multi-Teacher Knowledge Distillation (MTKD) framework. In the O-P strategy, we have already trained the models $\mathcal{M}_{O}$, $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, and $\mathcal{M}_{T_{L}}$. Building on this, we further train a student model $\mathcal{M}_{S}$. First, we initialize the student model $\mathcal{M}_{S}$ using the parameters from $\mathcal{M}_{O}$, and then use $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, and $\mathcal{M}_{T_{L}}$ as teacher models to perform KD. For each input image pair, we select the appropriate teacher model based on the image’s CAR to guide the student model. The process is shown in the green box at the bottom of Fig. 3. In this framework, when the student model $\mathcal{M}_{S}$ is supervised by the ground truth labels, it also receives the CM information corresponding to different CAR ranges, provided by the teacher models.
@@ V-B Benchmark Methods
To comprehensively verify the validity of our JL1-CD dataset and the superiority of the proposed O-P strategy and MTKD framework, we conduct extensive experiments on existing benchmark algorithms. The selected models, along with their corresponding backbones, parameter sizes, and computational complexities, are summarized in Table II. Based on the backbone architecture, the models are categorized into three groups: Alice Blue for CNN-based models, Light Cyan for Transformer-based models, and Lavender Blue for FM-based models, encompassing almost all mainstream architectures. FLOPs is calculated with an input image of size 512 × 512. As shown in the table, the selected models span a wide range of sizes, from lightweight models such as TinyCD and LightCDNet with less than 1MB to the latest SOTA model TTP, which exceeds 360MB. This wide range allows us to verify the universality of the O-P and MTKD methods across models with different backbones and scales.
@@ V-C Evaluation Metrics
It is important to note that in the toolboxes like MMSegmentation[53] and MMRotate [54], these metrics are computed based on the total number of pixels across all predicted label images. In this paper, however, to align with the competition requirements, we first calculate these metrics for each individual image and then compute the average of the results across all images to obtain the final outcome.
@@ V-D Implementation Details
All algorithms are trained and tested using the PyTorch-based OpenCD Toolbox [55]. To ensure a fair comparison and clearly assess the contributions of the Origin-Partition (O-P) strategy and the Multi-Teacher Knowledge Distillation (MTKD) framework to the performance improvement of the CD models, we adopt consistent settings for all models ($\mathcal{M}_{O}$, $\mathcal{M}_{T_{L}}$, $\mathcal{M}_{T_{M}}$, $\mathcal{M}_{T_{S}}$, and $\mathcal{M}_{S}$) across the various algorithms, as described below:
@@
When partitioning the dataset, we set the thresholds $th_{1}=0.05$ and $th_{2}=0.2$, which ensures a balanced distribution of samples across the partitions. The model is saved every 1k iterations (5 epochs for TTP), and the checkpoint with the highest mIoU value on the validation set is selected for testing.
@@ V-E Experimental Results
    Surprisingly, unlike their performance on the validation set, many algorithms under O-P perform worse than under MTKD, and in some cases, even worse than the original models (e.g., STANet, IFN, Changer-MiT-b1, etc.). However, the student models obtained through MTKD training can achieve significantly better performance than models trained under the original and O-P methods. This is mainly due to the inability of the original models to achieve accurate CAR on the test set during the coarse detection stage. This finding also indicates that our MTKD method can endow student models with superior capabilities compared to the teacher models, without increasing computation and time cost during inference.
@@
In remote sensing change detection, four main challenges are typically encountered: false alarms, missed detections, internal density, and boundary completeness. To visually assess the effectiveness of the O-P and MTKD strategies in addressing these challenges, we present the visual results of several algorithms in Fig. 7, where red indicates missed detections and blue indicates false alarms. The first five columns show performance in terms of missed detection. As depicted, the original algorithms struggle to detect small changes accurately, particularly for objects with subtle variations. However, after optimization with O-P and MTKD, the missed detection rate for small-scale changes is significantly reduced. Of course, the improvement is limited; for example, in the first and fourth columns, although MTKD detects more road changes, it still cannot fully capture the narrow variations. The sixth column displays the performance in terms of false alarms, where it is evident that the false alarm rate decreases progressively across the original, O-P, and MTKD methods. The last three columns illustrate the improvements in internal density and boundary completeness. We effectively reduce the occurrence of internal holes within the change areas, and the boundaries become more complete.
Results on Different CAR Partitions
We select one model from each of the CNN, Transformer, and FM architectures and evaluate their performance across different CAR ranges. Fig. 8 summarizes the mIoU results of HANet, ChangeFormer-MiT-b1, and TTP on both the validation set (first row) and the test set (second row). The images are sorted by CAR in ascending order and divided into five equal partitions. Each bar in the figure represents the lower and upper bounds of CAR for each partition, while the line graph indicates the mIoU across the different partitions. From the figure, it can be observed that the detection performance of O-P and MTKD may actually decrease for images with high CAR (e.g., ChangeFormer-MiT-b1 and TTP). However, O-P and MTKD significantly enhance model performance on images with low CAR. In the first partition of the test set (CAR range 0.02% to 0.88%), O-P improves the mIoU for the three algorithms by 8.15%, 3.21%, and 6.55%, respectively, while MTKD improves it by 4.42%, 3.36%, and 2.85%. These results demonstrate a significant improvement in the accuracy of detecting tiny changes, which is consistent with the findings in Fig. 7.
Comparison of Results on Change and No-Change Classes

### v2

@@ Abstract
Deep learning has achieved significant success in the field of remote sensing image change detection (CD), yet two major challenges remain: the scarcity of sub-meter, comprehensive open-source CD datasets, and the difficulty of achieving consistent and satisfactory detection results across images with varying change areas. To address these issues, we introduce the JL1-CD dataset, which consists of 5,000 pairs of 512 × 512 pixel images with a resolution of 0.5 to 0.75 meters. This all-inclusive dataset covers a wide range of human-induced and natural changes, including buildings, roads, hardened surfaces, woodlands, grasslands, croplands, water bodies, and photovoltaic panels, among others. Additionally, we propose a novel multi-teacher knowledge distillation (MTKD) framework that leverages the Origin-Partition (O-P) strategy to enhance CD performance. In the O-P strategy, we partition the training data based on the Change Area Ratio (CAR) to train separate models for small, medium, and large CAR values, alleviating the learning burden on each model and improving their performance within their respective partitions. Building upon this, our MTKD framework distills knowledge from multiple teacher models trained on different CAR partitions into a single student model,enabling the student model to achieve superior detection results across diverse CAR scenarios without incurring additional computational or time overhead during the inference phase. Experimental results on the JL1-CD and SYSU-CD datasets demonstrate that the MTKD framework significantly improves the performance of CD models with various network architectures and parameter sizes, achieving new state-of-the-art results. The JL1-CD dataset and code are available at https://github.com/circleLZY/MTKD-CD.
@@ I INTRODUCTION
In recent years, the rapid advancement of deep learning (DL) has revolutionized remote sensing CD, delivering substantial performance breakthroughs. DL-based CD methods generally involve three steps: (1) extracting change features from image pairs, (2) generating change maps based on the extracted features, and (3) predicting labels based on the feature maps. Convolutional Neural Networks (CNNs), which achieved remarkable success in image processing, were the first neural network architecture applied to remote sensing CD and remain widely optimized and utilized today [10, 11, 12]. With the introduction of Transformers, some studies have explored their application in CD tasks [13, 14]. More recently, Foundational Model (FM) has emerged as a novel paradigm, aiming to achieve multi-task and multi-domain generalization through large-scale pretraining [15].
However, DL-based CD methods generally face two major challenges: the scarcity of high-quality, high-resolution, all-inclusive CD datasets and limitations in handling highly dynamic change areas. Although numerous CD datasets have been constructed and proposed, they are often tailored to specific scenarios, which restricts the generalization capabilities of the algorithms. For instance, models trained on datasets focused on human-induced changes often fail to perform effectively when confronted with natural change scenarios. On the other hand, the learning capacity of these models is inherently limited. Most existing algorithms rely on a single-phase training approach, typically end-to-end training. While such training strategies can achieve satisfactory results within a constrained range of changes, the models’ performance significantly degrades when addressing scenarios with wide variations in change areas.
Figure 1: Timeline of the development of mainstream DL-based CD methods. The numbers “1-5” and “6-10” denote publication periods in 2024 (January-May and June-October, respectively).
@@ II-B DL-Based CD
Figure 2: Sample images from the JL1-CD dataset. Each row, from top to bottom, represents: the image at time 1, the image at time 2, and the ground truth label. Each column corresponds to different change types: (a) Decrease in woodland; (b) Building changes; (c) Conversion of cropland to greenhouses; (d) Road changes; (e) Waterbody changes; (f) Surface hardening; and (g) Photovoltaic panel construction
@@
Recently, foundation models have become a new training paradigm. There have been works utilizing remote sensing data to fine-tune pretrained models such as Vision Transformers (ViT)[29], Segment Anything Model (SAM)[30], and Contrastive Language-Image Pretraining (CLIP)[31], achieving higher performance in CD tasks[32, 33, 34, 35]. Li et al. [32] proposed the Bi-Temporal Adapter Network (BAN), a universal FM-based framework for CD, which enhances existing models with minimal additional parameters and achieves significant performance improvements. Chen et al. [34] introduced Time Travelling Pixels (TTP), a method that integrates latent knowledge from the SAM model into CD, overcoming domain shifts and spatio-temporal complexities, demonstrating SOTA results on the LEVIR-CD[22] dataset. Zheng et al. [35] developed AnyChange, a zero-shot CD model built on the SAM that utilizes bitemporal latent matching for training-free adaptation, setting a new SOTA on the SECOND[36] benchmark and achieving significant improvements in both unsupervised and supervised CD tasks.
@@ III JL1-CD Dataset
We collect the number of types of changes, number of image pairs, image size, and resolution information of mainstream CD datasets in Table I. The SZTAKI AirChange Benchmark [43] contains 12 pairs of 952 × 640 and one pair of 1,048 × 724 optical aerial images. It is one of the earliest and most commonly used CD datasets in early research. The DSIFN dataset [21] consists of 6 large bi-temporal image pairs from 6 cities in China, which are cropped into 394 sub-image pairs, each sized 512 × 512. The SECOND dataset [36] includes 4,662 pairs of aerial images collected from multiple platforms and sensors, covering cities such as Hangzhou, Chengdu, and Shanghai. Unlike other datasets that classify changes into only two categories (change and no change), SECOND provides detailed annotations for change types, including non-vegetated surfaces, trees, low vegetation, water, buildings, and playgrounds. However, the resolution of all or part of the images in these datasets does not reach the sub-meter level. WHU-CD [44], LEVIR-CD [22], and S2Looking [45] are very popular datasets that are specifically designed for monitoring building changes. These datasets predominantly include human-induced changes and lack natural change types. The CDD dataset[46] is derived from 7 pairs of 4,725 × 2,700 real-world seasonal change remote sensing images. The SYSU-CD dataset[47] contains 20,000 pairs of 0.5-meter aerial images captured in Hong Kong between 2007 and 2014. These datasets feature high resolution and diverse change types.
To provide a better benchmark for evaluating CD algorithms, we propose the JL1-CD dataset, a high-resolution, all-inclusive change detection dataset. JL1-CD includes 5,000 pairs of satellite images captured in China from early 2022 to the end of 2023, including Shandong, Ningxia, Anhui, Hebei, Hunan, and other regions. The images have sub-meter resolutions ranging from 0.5 to 0.75 meters and are sized 512 × 512 pixels. As shown in Fig. 2, the dataset covers various common human-induced and natural surface features, such as buildings, roads, hardened surfaces, woodlands, grasslands, croplands, water bodies and photovoltaic (PV) panels. It is worth noting that land cover changes induced by the construction of PV panels represent a significant change type in this dataset. As illustrated in Fig. 2 (g), these panels appear as black rectangles in the imagery due to their efficient sunlight absorption design. The resulting change regions exhibit diverse shapes and sizes, effectively enriching the diversity of change samples in our dataset. The original 5,000 image pairs are divided into 4,000 pairs for training and 1,000 pairs for testing, following an 80:20 split. The JL1-CD dataset will be made openly available for all research needs.
@@ IV-A O-P Strategy
The traditional training and testing approach for CD models is illustrated in the upper-left corner of Fig. 3 (green box) and the upper-right corner (pink box). For a given CD model $\mathcal{M}$, the input consists of a pair of images $(X_{1},X_{2})$, and the output is a change map (CM). For binary CD tasks with a single channel $c=1$, the predicted label $\hat{Y}$ is typically determined using a thresholding method:
@@
where $i$ denotes the pixel location in the image, “1” represents a change, and “0” represents no change (for visualization purpose, “1” will be mapped to a grayscale value of 255). The threshold $th$ is a predefined value. If $c=2$, the predicted label $\hat{Y}$ is generally determined via an element-wise comparison between the two layers of the change map, as follows:
@@
where $CM_{0}$ and $CM_{1}$ represent the first and second layers of the change map, respectively. If $c>2$, the CD model $\mathcal{M}$ not only predicts the locations of changes but also detects the types of changes, which is beyond the scope of this paper.
@@ IV-B MTKD Framework
By partitioning the training set, we effectively reduce the learning burden for each model. As a result, the Origin-Partition (O-P) strategy significantly enhances the performance of the CD algorithm. However, the O-P strategy still has two significant limitations: First, during inference, we are required to load four different models, and even disregarding data throughput, the time required is at least twice that of the original algorithm (often much more), significantly increasing both memory and time complexity. Second, since we first use $\mathcal{M}_{O}$ to obtain a rough estimate of the CAR, any inaccuracies in this estimation may lead to incorrect model selection in the subsequent steps, thereby reducing the accuracy of the final predictions. We are thus prompted to consider: is there a way to combine the capabilities of these four models into a single model?
To address this, we propose the Multi-Teacher Knowledge Distillation (MTKD) framework. In the O-P strategy, we have already trained the models $\mathcal{M}_{O}$, $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, and $\mathcal{M}_{T_{L}}$. Building on this, we further train a student model $\mathcal{M}_{S}$. First, we initialize the student model $\mathcal{M}_{S}$ using the parameters from $\mathcal{M}_{O}$, and then use $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, and $\mathcal{M}_{T_{L}}$ as teacher models to perform KD. For each input image pair, we select the appropriate teacher model based on the image’s CAR to guide the student model. The process is shown in the green box at the bottom of Fig. 3. In this framework, the student model $\mathcal{M}_{S}$ is simultaneously supervised by the ground truth labels and the CM information from the teacher models across different CAR partitions.
@@ V-B Benchmark Methods
To comprehensively verify the validity of our JL1-CD dataset and the superiority of the proposed O-P strategy and MTKD framework, we conduct extensive experiments on existing benchmark algorithms. The selected models, along with their corresponding backbones, parameter sizes, and computational complexities, are summarized in Table II. Based on the backbone architecture, the models are categorized into three groups: Alice Blue for CNN-based models, Light Cyan for Transformer-based models, and Lavender Blue for FM-based models, encompassing almost all mainstream architectures. FLOPs is calculated with an input image of size 512 × 512. As shown in the table, the selected models span a wide range of sizes, from lightweight models with fewer than 1M parameters, such as TinyCD and LightCDNet, to the latest SOTA model TTP with over 360M parameters. The inclusion of these lightweight models is crucial, as improving their detection performance without increasing inference resource consumption is of significant importance for deployment in edge devices and scenarios requiring real-time processing. This wide range of model sizes also allows us to verify the universality of the O-P and MTKD methods across models with different backbones and scales.
@@ V-C Evaluation Metrics
It is important to note that in the MMSegmentation[50] toolbox, these metrics are computed based on the total number of pixels across all predicted label images. In this paper, however, to align with the competition requirements, we first calculate these metrics for each individual image and then compute the average of the results across all images to obtain the final outcome.
@@ V-D Implementation Details
All algorithms are trained and tested using the PyTorch-based OpenCD Toolbox [51]. To ensure fair comparisons and to clearly assess the contributions of the Origin-Partition (O-P) strategy and the Multi-Teacher Knowledge Distillation (MTKD) framework to the performance improvement of the CD models, we adopt consistent parameter settings and hardware conditions for all models ($\mathcal{M}_{O}$, $\mathcal{M}_{T_{L}}$, $\mathcal{M}_{T_{M}}$, $\mathcal{M}_{T_{S}}$, and $\mathcal{M}_{S}$) across the various algorithms, as described below:
@@
When partitioning the dataset, we set the thresholds $th_{1}=0.05$ and $th_{2}=0.2$ by empirical, which ensures a balanced distribution of samples across the partitions. The model is saved every 1k iterations (5 epochs for TTP), and the checkpoint with the highest mIoU value on the validation set is selected for testing.
@@ V-E Experimental Results
    It is understandable that many models, such as SNUNet, BIT, and Changer-r18, perform better under the O-P strategy than under MTKD, as for well-trained models, the performance of teacher models should naturally surpass that of the student model. In the O-P strategy, the output from one of the teacher models is used as the final result, which allows its performance on the test set to exceed that of MTKD. However, it is equally noteworthy that several algorithms, such as STANet, IFN, and Changer-MiT-b1, perform worse under the O-P strategy than under MTKD, and in some cases, even worse than their original models. This is primarily because the original models fail to accurately estimate the CAR during the coarse detection stage on the test set. As a result, the subsequent selection of the wrong teacher model for further testing leads to a degradation in detection performance. This behavior highlights the potential shortcomings of the O-P strategy in certain scenarios. This finding also demonstrates that our MTKD method can enhance student models’ capabilities beyond those of the teacher models trained on partitioned data, without increasing computational or time costs during inference.
@@
    The lightweight model TinyCD demonstrates competitive performance on the JL1-CD dataset, comparable to that of IFN (which has over 100 times more parameters). When trained within the MTKD framework, TinyCD’s mIoU and mFscore improve by 1.51% and 1.52%, respectively. This is crucial for resource-constrained edge device scenarios. However, while the O-P strategy significantly improves the performance of another lightweight model LightCDNet (with improvements of 3.49% in mIoU and 3.70% in mFscore), the MTKD framework leads to a performance degradation. This suggests that further improvements are needed in our approach when optimizing performance for lightweight models.
  4. 4)
@@
In remote sensing CD, four main challenges are typically encountered: false alarms, missed detections, internal density, and boundary completeness. To visually assess the effectiveness of the O-P and MTKD strategies in addressing these challenges, we present the visual results of several algorithms in Fig. 7, where red indicates missed detections and blue indicates false alarms. In our experiments, we identify missed detections and false alarms as the most common issues, which are illustrated in the first six columns of Fig. 7. In Fig. 7(a), the original BAN algorithm exhibits many false alarms when detecting large CAR changes caused by PV panels and nearly fails to detect a newly constructed road on the right side of the image. After O-P optimization, false alarms related to PV panels are significantly reduced, but the road detection performance deteriorates. In contrast, MTKD achieves the best balance in detecting both small and large CAR changes, reducing false alarms related to PV panels while minimizing missed detections of the road. Fig. 7(c) similarly shows the results for detecting narrow roads. After MTKD optimization, SNUNet’s detection of the changed roads significantly improves in terms of completeness and connectivity, though some missed detections remain. his indicates that detecting changes caused by narrow roads remains a challenging task that requires further research. Our method not only improves the detection of large changes caused by buildings (Fig. 7(b)) and water body expansions (Fig. 7(e)) but also demonstrates notable advantages in detecting small object changes. Fig. 7(b) and Fig. 7(f) present results for small CAR scenarios. Whether qualitatively, as observed in the images, or quantitatively, as shown in the subsequent experiments, detecting small-scale changes remains a major challenge for all current CD algorithms. However, after optimization with O-P and MTKD, false alarms and missed detections in small-scale change scenarios are significantly improved. Of course, the improvement is limited; for instance, in the second and fourth columns, although MTKD detects more changes, it still cannot fully capture the tiny variations. This indicates that further improvements are needed in detecting small CAR changes. The last three columns illustrate the improvements in internal density and boundary completeness. We effectively reduce the occurrence of internal holes within the change areas, and the boundaries become more complete.
@@
Results on Different CAR Partitions
We select one model from each of the CNN, Transformer, and FM architectures and evaluate their performance across different CAR ranges. Fig. 8 summarizes the mIoU results of HANet, ChangeFormer-MiT-b1, and TTP on both the validation set (first row) and the test set (second row). The images are sorted by CAR in ascending order and divided into five equal partitions. Each bar in the figure represents the lower and upper bounds of CAR for each partition, while the line graph indicates the mIoU across the different partitions. From the figure, it can be observed that the detection performance of O-P and MTKD may actually decrease for images with high CAR (e.g., ChangeFormer-MiT-b1 and TTP). However, O-P and MTKD significantly enhance model performance on images with low CAR. In the first partition of the test set (CAR range 0.02% to 0.88%), O-P improves the mIoU for the three algorithms by 8.15%, 3.21%, and 6.55%, respectively, while MTKD improves it by 4.42%, 3.36%, and 2.85%. These results show that, due to the wide CAR range of the JL1-CD dataset, directly training existing algorithms using their original models can result in errors when detecting small changes. However, by partitioning small CAR regions in the O-P strategy and transferring the knowledge from the corresponding teacher model to the student model in MTKD, our methods achieve significant improvements in the accuracy of detecting tiny changes, which is consistent with the findings in Fig. 7.
Comparison of Results on Change and No-Change Classes